x-y=2
x^2+y^2=244

Answer by ikleyn(52781)

.
x-y=2
x^2+y^2=244
---------------------

Express  x  from the first equation,  x = y+2,  and then substitute it into the second equation.

You will get a quadratic equation for the single unknown  y.

Solve it.

Good luck!